A certain group of test subjects had pulse rates with a mean of

82.3

beats per minute and a standard deviation of

10.2

beats per minute. Use the range rule of thumb to identify the limits separating values that are significantly low or significantly high. Is a pulse rate of

112.7

beats per minute significantly low or significantly​ high?

Significantly low values are

nothing

beats per minute or lower.

​(Type an integer or a decimal. Do not​ round.)

Significantly high values are

nothing

beats per minute or higher.

​(Type an integer or a decimal. Do not​ round.)

Is a pulse rate of

112.7

beats per minute significantly low or significantly​ high?
Answer #1

ANSWER:

Given THAT

= 82.3, = 10.2

Here

- 2 = 82.3 - 2 * 10.2 = 61.9 and + 2 = 82.3 + 2 * 10.2 = 102.7

Thus 61.9 is significantly low and

102.7 is significantly high

Given pulse rate is 112.7 beats per minutes which is greater than upper limit.

Therefore, pulse rate of 112.7 beats per minutes is significantly high.